# Quantum metric-based optical selection rules

### v1

@@ Abstract
Introduction.—In atoms, optical selection rules govern allowed transitions based on orbital angular momentum and the symmetry of atomic orbitals [1]. Inspired by the close connection between optical selection rules and orbital angular momentum in atoms, the discovery that orbital angular momentum in solids is closely related to Berry curvature led to the emergence of Berry curvature-based optical selection rules for circularly polarized light [2, 3, 4, 5, 6, 7], as schematically illustrated in Fig. 1(b). These rules are well established and widely studied, particularly in graphene and transition metal dichalcogenides [5, 6, 7]. However, both the Berry curvature and the quantum metric—representing the imaginary and real parts of quantum geometry, respectively—have been shown to qualitatively influence optical transitions [8, 9, 10, 11, 12, 13, 14, 15, 16]. Consequently, the quantum metric is also expected to give rise to optical selection rules. Moreover, as both linear and circular polarization are fundamental and equally significant polarization states of light, optical selection rules should extend beyond circularly polarized light to include linearly polarized light, considering that some works [17, 18, 19, 20] have noted certain valleys can exhibit complete linear polarization selection.
@@
In this work, we establish quantum metric-based optical selection rules for linearly polarized light, linking the quantum metric directly to oscillator strength for linear polarization of light, as schematically illustrated in Fig. 1(a). For a valley with mirror symmetry, the electrons in the valley couple exclusively to linear polarization aligned parallel/perpendicular to its mirror plane. When two valleys exhibit mutually perpendicular mirror planes and are connected by mirror symmetry or four-fold rotational symmetry, we demonstrate valley-contrasted selection rules: orthogonal linear polarizations selectively excite distinct valleys. We validate our theory in an altermagnet (AM) model and the Kane-Mele (KM) model, and further probe its realization in monolayer $d$-wave altermagnet $\mathrm{V_{2}SeSO}$ via density functional theory (DFT). This work extends the geometric origin of optical selection rules beyond the Berry curvature paradigm.
@@
FIG. 1: Schematics of the quantum metric-based optical selection rules for linearly polarized light (a) and the Berry curvature-based optical selection rules for circularly polarized light (b). (a) The two two-band valleys are located at mirror-invariant lines or rotation axes of in-plane two-fold rotational symmetry and related by $M$ symmetry (e.g., mirror symmetry or four-fold rotational symmetry) with non-zero quantum metric $g^{xy}(\boldsymbol{k})=-g^{xy}(M\boldsymbol{k})$ and vanishing Berry curvature. For two orthogonal linearly polarized lights along specific directions, one (another) light exclusively excites electrons at one (the other) valley. (b) The two two-band valleys are located at $n$-fold rotation axes ($n\geq 3$) and related by $T$ symmetry (e.g., time-reversal symmetry or mirror symmetry) with non-zero Berry curvature $\Omega^{xy}(\boldsymbol{k})=-\Omega^{xy}(T\boldsymbol{k})$ and vanishing quantum metric. The left (right) circularly polarized light exclusively excites electrons at one (the other) valley.
General theory.—First, we revisit the Berry curvature-based optical selection rules for circularly polarized light with the details in Supplementary Materials (SM) [21]. One can assume the light propagates along the $z$-direction. There is a link between the Berry curvature $\Omega_{nm}^{xy}$ and the oscillator strengths for left- and right-circularly polarized light $f_{nm}^{(\hat{\boldsymbol{x}}\pm i\hat{\boldsymbol{y}})}(\boldsymbol{k})$, i.e., $\Omega^{xy}_{nm}(\boldsymbol{k})=-\hbar(f_{nm}^{(\hat{\boldsymbol{x}}+i\hat{\boldsymbol{y}})}(\boldsymbol{k})-f_{nm}^{(\hat{\boldsymbol{x}}-i\hat{\boldsymbol{y}})}(\boldsymbol{k}))/{4m_{e}\omega_{nm,\boldsymbol{k}}}$, where $m_{e}$ is the electron mass and $\hbar\omega_{nm,\boldsymbol{k}}\equiv\hbar\omega_{n,\boldsymbol{k}}-\hbar\omega_{m,\boldsymbol{k}}$ is the energy difference between the $m$th and $n$th bands. The degree of circular polarization $\eta_{nm}(\boldsymbol{k})$ for the optical transition between states $n$ and $m$ is given by $\eta_{nm}(\boldsymbol{k})=-\Omega^{xy}_{nm}(\boldsymbol{k})/(g^{xx}_{nm}(\boldsymbol{k})+g^{yy}_{nm}(\boldsymbol{k}))$. A complete circular polarization selectivity emerges with $\eta(\boldsymbol{k})=\pm 1$ at a valley with $n$-fold ($n\geq 3$) rotational symmetry and non-zero Berry curvature. The left or right circularly polarized light couples exclusively to the optical transitions at the valley. When two such valleys are connected by time-reversal symmetry or mirror symmetry, there are the valley-contrasted optical selection rules, i.e., one (the other) valley selectively couples with left (right) circularly polarized light, as shown in Fig. 1(b).
@@
Now, we derive the quantum metric-based optical selection rules for linearly polarized light. The matrix elements of the quantum metric tensor $\boldsymbol{g}(\boldsymbol{k})$ can be given by [8, 9, 10, 11]
@@
The matrix elements of the quantum metric tensor, i.e., $g^{xy}_{nm}(\boldsymbol{k})$, can be rewritten in terms of the oscillator strength for the transition between states $n$ and $m$ as (SM [21])
$g^{xy}_{nm}(\boldsymbol{k})=\dfrac{\hbar(f_{nm}^{(\hat{\boldsymbol{x}}+\hat{\boldsymbol{y}})}(\boldsymbol{k})-f_{nm}^{(\hat{\boldsymbol{x}}-\hat{\boldsymbol{y}})}(\boldsymbol{k}))}{8m_{e}\omega_{nm,\boldsymbol{k}}}.$ (2)
$f_{nm}^{(\hat{\boldsymbol{x}}+\hat{\boldsymbol{y}})}(\boldsymbol{k})$ and $f_{nm}^{(\hat{\boldsymbol{x}}-\hat{\boldsymbol{y}})}(\boldsymbol{k})$ are the $k$-resolved oscillator strengths, i.e., the optical absorption strengths, for linear polarization along the $\hat{\boldsymbol{x}}+\hat{\boldsymbol{y}}$ and $\hat{\boldsymbol{x}}-\hat{\boldsymbol{y}}$, respectively, and read [22, 23, 24]
@@
We now consider the scenario of optical transitions involving two bands, where $\eta(\boldsymbol{k})$ can be simplified to
@@
The quantum metric tensor is positive semi-definite [8, 9, 25, 26, 27], i.e., $g_{vc}^{xx}(\boldsymbol{k})g_{vc}^{yy}(\boldsymbol{k})-g_{vc}^{xy}(\boldsymbol{k})g_{vc}^{yx}(\boldsymbol{k})\geq 0$, and is symmetric under swapping of $x$ and $y$, i.e., $g_{vc}^{xy}(\boldsymbol{k})=g_{vc}^{yx}(\boldsymbol{k})$. Therefore, $-1\leq\eta(\boldsymbol{k})\leq 1$ and the inequality is saturated when $g^{xy}_{vc}(\boldsymbol{k})=g^{xx}_{vc}(\boldsymbol{k})=g^{yy}_{vc}(\boldsymbol{k})$ or $-g^{xy}_{vc}(\boldsymbol{k})=g^{xx}_{vc}(\boldsymbol{k})=g^{yy}_{vc}(\boldsymbol{k})$. We find that $\eta(\boldsymbol{k})$ takes the values of ±$\pm 1$ when $\boldsymbol{k}$ lies on the mirror-invariant planes. Specifically, the mirror symmetry $M_{xy}:(x,y)\to(-y,-x)$ imposes $\langle m|v^{x}_{\boldsymbol{k}}|n\rangle=\langle M_{xy}m|(-v^{y}_{\boldsymbol{k}})|M_{xy}n\rangle=-e^{i(\theta_{n}-\theta_{m})}\langle m|v^{y}_{\boldsymbol{k}}|n\rangle$, where $e^{i\theta_{n}}$ is the mirror eigenvalue for the state $|n\rangle$, and thus
@@
Substituting Eq. (6) and Eq. (7) separately into Eq. (5), we find that both $M_{xy}$ and $M_{x,-y}$ can ensure $\eta=\pm 1$. Therefore, the specific mirror symmetry $M$ $(M=M_{xy},M_{x,-y})$ impose a selection rule that the linearly polarized light polarized in $\hat{\boldsymbol{x}}+\hat{\boldsymbol{y}}$ or $\hat{\boldsymbol{x}}-\hat{\boldsymbol{y}}$ is exclusively allowed in optical transitions at the mirror invariant plane of $M$, i.e., $\boldsymbol{k}=M\boldsymbol{k}$. As a result, two-band valleys with mirror symmetry and nonzero quantum metric exhibit complete linear polarization selectivity. Note that if $\boldsymbol{k}$ simultaneously lies in the mirror invariant planes of $M_{xy}$ and $M_{x,-y}$, the complete linear polarization selectivity can still emerge when the mirror eigenvalues of the conduction and valence bands are identical for one mirror symmetry but opposite for the other, as derived from Eqs. (6)(7).
However, the complete linear polarization selectivity cannot be exhibited in many-band valleys. The reason is that the inequality arising from the positive-definiteness of the quantum metric tensor holds with equality only in the two-band case [25, 26, 27], and thus $\eta(\boldsymbol{k})$ cannot be ±$\pm 1$. The complete circular polarization selectivity in Berry curvature-based optical selection rules can also only be achieved in two-band valleys due to the positive semi-definiteness of the quantum geometry tensor (SM [21]).
@@
where $\xi=1$ at X and $\xi=-1$ at Y. The velocity operators are given by $v^{x}=\xi v^{y}=-\sqrt{2}t\sigma_{x}/\hbar$. According to Eqs. (1)(5), we have $g^{xy}_{vc}=g^{xx}_{vc}=g^{yy}_{vc}$ at X and $-g^{xy}_{vc}=g^{xx}_{vc}=g^{yy}_{vc}$ at Y, and thus $\eta(\mathrm{X})=-\eta(\mathrm{Y})=1$. Therefore, optical transitions at the X valley exclusively admit $\hat{\boldsymbol{x}}+\hat{\boldsymbol{y}}$-polarized light, while the Y valley selectively responds to $\hat{\boldsymbol{x}}-\hat{\boldsymbol{y}}$ polarization. The numerically calculated oscillator strengths and degree of polarization are shown in Fig. 2(d1), and $\eta(\boldsymbol{k})$ are exactly ±$\pm 1$ on the mirror-invariant lines.
@@
For linearly polarized light with $\hat{\boldsymbol{x}}+\sqrt{3}\hat{\boldsymbol{y}}$ and $\sqrt{3}\hat{\boldsymbol{x}}-\hat{\boldsymbol{y}}$ polarizations at M, the degree of polarization is given by $\eta(\mathrm{M})=(t^{2}-12t^{2}_{\mathrm{SOC}})/(t^{2}+12t^{2}_{\mathrm{SOC}})$. Similarly, the degree of polarization for linearly polarized light with $\hat{\boldsymbol{x}}-\sqrt{3}\hat{\boldsymbol{y}}$ and $\sqrt{3}\hat{\boldsymbol{x}}+\hat{\boldsymbol{y}}$ at M are given by $\eta(\mathrm{M}^{\prime})=-(t^{2}-12t^{2}_{\mathrm{SOC}})/(t^{2}+12t^{2}_{\mathrm{SOC}})$. $\eta(\mathrm{M})$ ($\eta(\mathrm{M^{\prime}})$) can take the value of 1 (-1) when $t_{\mathrm{SOC}}=0$. See SM [21] for more details.
@@
The monolayer $\mathrm{V_{2}SeSO}$ has two valleys at X and Y, and the two valleys are locked to opposite spins, as shown in Fig. 3(c). The two valleys are related by the mirror symmetries $M_{x}$ and $M_{y}$. Therefore, the degree of linear polarization $\eta(\boldsymbol{k})$ is opposite around the two valleys as shown in Fig. 3(d). Moreover, both the mirror-invariant lines of $M_{xy}$ and $M_{x,-y}$ pass through the two valleys. For the X valley, the conduction and valence bands have identical $M_{x,-y}$ eigenvalues but have different $M_{xy}$ eigenvalues. For the Y valley, the conduction and valence bands have identical $M_{xy}$ eigenvalues but have different $M_{x,-y}$ eigenvalues. From Eq. (6) and Eq. (7), we have $g^{xy}_{vc}(\boldsymbol{k})=g^{xx}_{vc}(\boldsymbol{k})=g^{yy}_{vc}(\boldsymbol{k})$ for valley at X and $-g^{xy}_{vc}(\boldsymbol{k})=g^{xx}_{vc}(\boldsymbol{k})=g^{yy}_{vc}(\boldsymbol{k})$ for valley at Y. Therefore, $\eta(\mathrm{X})=-\eta(\mathrm{Y})=1$ according to Eq. (5). The $\hat{\boldsymbol{x}}+\hat{\boldsymbol{y}}$-polarized light is exclusively allowed at X valley and the $\hat{\boldsymbol{x}}-\hat{\boldsymbol{y}}$-polarized light is exclusively allowed at Y valley. As shown in Fig. 3(d), the overwhelming majority of optical transitions around the valleys preserve complete linear polarization selectivity.
@@
The AM has garnered significant interest due to its null net magnetization and unique spin-splitting bands that facilitate spin-valley locking [28, 41], and thus is a promising platform for spintronics, valleytronics, and optoelectronics applications. Achieving complete circular polarization selectivity in Berry curvature-based optical selection rules usually requires valleys with three-fold rotational symmetry. However, valleys in two-dimensional AMs with three-fold or higher rotational symmetry are generally spin-degenerate [28, 41]. The valleys with mirror or mirror-like symmetry in our proposed quantum metric-based optical selection rules can be spin-splitting. Therefore, AMs are a good playground for our proposed quantum metric-based optical selection rules.

### v2

@@ Abstract
Introduction.—In atoms, optical selection rules govern allowed transitions based on atomic orbitals’ orbital angular momentum and symmetries [1]. Inspired by the close connection between optical selection rules and orbital angular momentum in atoms, the discovery that orbital angular momentum in solids is closely related to Berry curvature led to the emergence of Berry curvature-based optical selection rules for circularly polarized light [2, 3, 4, 5, 6, 7], as schematically illustrated in Fig. 1(b). These rules are well established and widely studied, particularly in graphene and transition metal dichalcogenides [5, 6, 7]. However, both the Berry curvature and the quantum metric—representing the imaginary and real parts of quantum geometry, respectively—have been shown to qualitatively influence optical transitions [8, 9, 10, 11, 12, 13, 14, 15, 16]. Consequently, the quantum metric is also expected to give rise to optical selection rules. Moreover, as both linear and circular polarization are fundamental and equally significant polarization states of light, optical selection rules should extend beyond circularly polarized light to include linearly polarized light, considering that some works [17, 18, 19, 20, 21, 22] have noted certain valleys can exhibit complete linear polarization selection.
@@
In this work, we establish quantum metric-based optical selection rules for linearly polarized light, linking the quantum metric directly to oscillator strength for linear polarization of light, as schematically illustrated in Fig. 1(a). For a valley with mirror or mirror-like symmetry, the electrons in the valley couple exclusively to linear polarization aligned parallel/perpendicular to its mirror plane. When two valleys exhibit mutually perpendicular mirror planes and are connected by symmetry like mirror symmetry and four-fold rotational symmetry, we demonstrate valley-contrasted selection rules: orthogonal linear polarizations selectively excite distinct valleys. We validate our theory in an altermagnet (AM) model and the Kane-Mele (KM) model, and further probe its realization in monolayer $d$-wave altermagnet $\mathrm{V_{2}SeSO}$ via density functional theory (DFT). This work extends the geometric origin of optical selection rules beyond the Berry curvature paradigm.
@@
FIG. 1: Schematics of the quantum metric-based optical selection rules for linearly polarized light (a) and the Berry curvature-based optical selection rules for circularly polarized light (b). (a) The two valleys are located at mirror-invariant lines or rotation axes of in-plane two-fold rotational symmetry and related by $M$ symmetry (e.g., mirror symmetry or four-fold rotational symmetry) with non-zero quantum metric $g^{xy}(\boldsymbol{k})=-g^{xy}(M\boldsymbol{k})$ and vanishing Berry curvature. For two orthogonal linearly polarized lights along specific directions, one (another) light exclusively excites electrons at one (the other) valley. (b) The two valleys are located at $n$-fold rotation axes ($n\geq 3$) and related by $T$ symmetry (e.g., time-reversal symmetry or mirror symmetry) with non-zero Berry curvature $\Omega^{xy}(\boldsymbol{k})=-\Omega^{xy}(T\boldsymbol{k})$ and vanishing quantum metric. The left (right) circularly polarized light exclusively excites electrons at one (the other) valley.
@@
General theory.—Here, we derive the quantum metric-based optical selection rules for linearly polarized light. The matrix elements of the quantum metric tensor $\boldsymbol{g}(\boldsymbol{k})$ can be given by [8, 9, 10, 11]
@@
Without loss of generality, one can assume the light propagates along the $z$-direction. The matrix elements of the quantum metric tensor, i.e., $g^{xy}_{nm}(\boldsymbol{k})$, can be rewritten in terms of the oscillator strength for the transition between states $n$ and $m$ as (Supplementary Materials [23] Sec. I )
$g^{xy}_{nm}(\boldsymbol{k})=\dfrac{\hbar(f_{nm}^{(\hat{\boldsymbol{x}}+\hat{\boldsymbol{y}})}(\boldsymbol{k})-f_{nm}^{(\hat{\boldsymbol{x}}-\hat{\boldsymbol{y}})}(\boldsymbol{k}))}{8m_{e}\omega_{nm,\boldsymbol{k}}},$ (2)
where $m_{e}$ is the electron mass and $\hbar\omega_{nm,\boldsymbol{k}}\equiv\hbar\omega_{n,\boldsymbol{k}}-\hbar\omega_{m,\boldsymbol{k}}$ is the energy difference between the $m$th and $n$th bands. $f_{nm}^{(\hat{\boldsymbol{x}}+\hat{\boldsymbol{y}})}(\boldsymbol{k})$ and $f_{nm}^{(\hat{\boldsymbol{x}}-\hat{\boldsymbol{y}})}(\boldsymbol{k})$ are the $k$-resolved oscillator strengths, i.e., the optical absorption strengths, for linear polarization along the $\hat{\boldsymbol{x}}+\hat{\boldsymbol{y}}$ and $\hat{\boldsymbol{x}}-\hat{\boldsymbol{y}}$, respectively, and read [24, 25, 26]
@@
We now consider the scenario of optical transitions involving two energy levels, labeled as $E_{c}$ and $E_{v}$, that are non-degenerate. $\eta(\boldsymbol{k})$ can be simplified to
@@
The quantum metric tensor is positive semi-definite [8, 9, 27, 28, 29], i.e., $g_{vc}^{xx}(\boldsymbol{k})g_{vc}^{yy}(\boldsymbol{k})-g_{vc}^{xy}(\boldsymbol{k})g_{vc}^{yx}(\boldsymbol{k})\geq 0$, and is symmetric under swapping of $x$ and $y$, i.e., $g_{vc}^{xy}(\boldsymbol{k})=g_{vc}^{yx}(\boldsymbol{k})$. Therefore, $1\leq\eta(\boldsymbol{k})\leq 1$ and we find that $\eta(\boldsymbol{k})$ takes the values of $\pm 1$ when $\boldsymbol{k}$ lies on the mirror-invariant planes, with $g^{xy}_{vc}(\boldsymbol{k})=g^{xx}_{vc}(\boldsymbol{k})=g^{yy}_{vc}(\boldsymbol{k})$ or $g^{xy}_{vc}(\boldsymbol{k})=g^{xx}_{vc}(\boldsymbol{k})=g^{yy}_{vc}(\boldsymbol{k})$. Specifically, the mirror symmetry $M_{xy}:(x,y)\to(-y,-x)$ imposes $\langle m|v^{x}_{\boldsymbol{k}}|n\rangle=\langle M_{xy}m|(-v^{y}_{\boldsymbol{k}})|M_{xy}n\rangle=-e^{i(\theta_{n}-\theta_{m})}\langle m|v^{y}_{\boldsymbol{k}}|n\rangle$, where $e^{i\theta_{n}}$ is the mirror eigenvalue for the state $|n\rangle$, and thus
@@
Substituting Eq. (6) and Eq. (7) separately into Eq. (5), we find that both $M_{xy}$ and $M_{x,-y}$ can ensure $\eta=\pm 1$. Therefore, the specific mirror symmetry $M$ $(M=M_{xy},M_{x,-y})$ impose a selection rule that the linearly polarized light polarized in $\hat{\boldsymbol{x}}+\hat{\boldsymbol{y}}$ or $\hat{\boldsymbol{x}}-\hat{\boldsymbol{y}}$ is exclusively allowed in optical transitions at the mirror invariant plane of $M$, i.e., $\boldsymbol{k}=M\boldsymbol{k}$. As a result, a valley exhibits complete linear polarization selectivity if the optical transition at the valley involves nondegenerate two energy levels and the valley has mirror symmetry and nonzero quantum metric (SM [23] Secs. I and II). Note that if $\boldsymbol{k}$ simultaneously lies in the mirror invariant planes of $M_{xy}$ and $M_{x,-y}$, the complete linear polarization selectivity can still emerge when the mirror eigenvalues of the conduction and valence bands are identical for one mirror symmetry but opposite for the other, as derived from Eqs. (6) and (7).
Here we revisit the Berry curvature-based optical selection rules for circularly polarized light with the details in SM [23] Sec. I. There is a link between the Berry curvature $\Omega_{nm}^{xy}$ and the oscillator strengths for left- and right-circularly polarized light $f_{nm}^{(\hat{\boldsymbol{x}}\pm i\hat{\boldsymbol{y}})}(\boldsymbol{k})$, i.e., $\Omega^{xy}_{nm}(\boldsymbol{k})=-\hbar(f_{nm}^{(\hat{\boldsymbol{x}}+i\hat{\boldsymbol{y}})}(\boldsymbol{k})-f_{nm}^{(\hat{\boldsymbol{x}}-i\hat{\boldsymbol{y}})}(\boldsymbol{k}))/{4m_{e}\omega_{nm,\boldsymbol{k}}}$. The degree of circular polarization $\eta_{nm}(\boldsymbol{k})$ for the optical transition between states $n$ and $m$ is given by $\eta_{nm}(\boldsymbol{k})=-\Omega^{xy}_{nm}(\boldsymbol{k})/(g^{xx}_{nm}(\boldsymbol{k})+g^{yy}_{nm}(\boldsymbol{k}))$. A complete circular polarization selectivity emerges with $\eta(\boldsymbol{k})=\pm 1$ at a valley with $n$-fold ($n\geq 3$) rotational symmetry, non-zero Berry curvature, and nondegenerate two energy levels (SM [23] Sec. I D). The left or right circularly polarized light couples exclusively to the optical transitions at a certain valley. When two such valleys are connected by time-reversal symmetry or mirror symmetry, there are the valley-contrasted optical selection rules, i.e., one (the other) valley selectively couples with left (right) circularly polarized light, as shown in Fig. 1(b).
The complete linear polarization selectivity cannot be exhibited in valleys where the optical transitions involve two energy levels that are degenerate. The reason is that the inequality arising from the positive semidefiniteness of the quantum metric tensor holds with equality only when the optical transitions involve nondegenerate two energy levels [27, 28, 29], and thus $\eta(\boldsymbol{k})$ cannot be $\pm 1$. The complete circular polarization selectivity in Berry curvature-based optical selection rules also can not be achieved in valleys with degenerate two energy levels due to the positive semidefiniteness of the quantum geometry tensor (SM [23] Sec. I D). However, optical transitions of a specific frequency generally involve nondegenerate two energy levels, so complete polarization selectivity can often be achieved.
@@
where $\xi=1$ at X and $\xi=-1$ at Y. The velocity operators are given by $v^{x}=\xi v^{y}=-\sqrt{2}t\sigma_{x}/\hbar$. According to Eqs. (1)(5), we have $g^{xy}_{vc}=g^{xx}_{vc}=g^{yy}_{vc}$ at X and $g^{xy}_{vc}=g^{xx}_{vc}=g^{yy}_{vc}$ at Y, and thus $\eta(\mathrm{X})=-\eta(\mathrm{Y})=1$. Therefore, optical transitions at the X valley exclusively admit $\hat{\boldsymbol{x}}+\hat{\boldsymbol{y}}-$polarized light, while the Y valley selectively responds to $\hat{\boldsymbol{x}}-\hat{\boldsymbol{y}}$ polarization. The numerically calculated oscillator strengths and degree of polarization are shown in Fig. 2(d1), and $\eta(\boldsymbol{k})$ are exactly $\pm 1$ on the mirror-invariant lines.
@@
For linearly polarized light with $\hat{\boldsymbol{x}}+\sqrt{3}\hat{\boldsymbol{y}}$ and $\sqrt{3}\hat{\boldsymbol{x}}-\hat{\boldsymbol{y}}$ polarizations at M, the degree of polarization is given by $\eta(\mathrm{M})=(t^{2}-12t^{2}_{\mathrm{SOC}})/(t^{2}+12t^{2}_{\mathrm{SOC}})$. Similarly, the degree of polarization for linearly polarized light with $\hat{\boldsymbol{x}}-\sqrt{3}\hat{\boldsymbol{y}}$ and $\sqrt{3}\hat{\boldsymbol{x}}+\hat{\boldsymbol{y}}$ at M are given by $\eta(\mathrm{M}^{\prime})=-(t^{2}-12t^{2}_{\mathrm{SOC}})/(t^{2}+12t^{2}_{\mathrm{SOC}})$. $\eta(\mathrm{M})$ ($\eta(\mathrm{M^{\prime}})$) can take the value of 1 (-1) when $t_{\mathrm{SOC}}=0$. See SM [23] Sec. III for more details.
@@
The monolayer $\mathrm{V_{2}SeSO}$ has two valleys at X and Y, and the two valleys are locked to opposite spins, as shown in Fig. 3(c). The two valleys are related by the mirror symmetries $M_{x}$ and $M_{y}$. Therefore, the degree of linear polarization $\eta(\boldsymbol{k})$ is opposite around the two valleys as shown in Fig. 3(d). Moreover, both the mirror-invariant lines of $M_{xy}$ and $M_{x,-y}$ pass through the two valleys. For the X valley, the conduction and valence bands have identical $M_{x,-y}$ eigenvalues but have different $M_{xy}$ eigenvalues. For the Y valley, the conduction and valence bands have identical $M_{xy}$ eigenvalues but have different $M_{x,-y}$ eigenvalues. From Eq. (6) and Eq. (7), we have $g^{xy}_{vc}(\boldsymbol{k})=g^{xx}_{vc}(\boldsymbol{k})=g^{yy}_{vc}(\boldsymbol{k})$ for valley at X and $g^{xy}_{vc}(\boldsymbol{k})=g^{xx}_{vc}(\boldsymbol{k})=g^{yy}_{vc}(\boldsymbol{k})$ for valley at Y. Therefore, $\eta(\mathrm{X})=-\eta(\mathrm{Y})=1$ according to Eq. (5). The $\hat{\boldsymbol{x}}+\hat{\boldsymbol{y}}-$polarized light is exclusively allowed at X valley and the $\hat{\boldsymbol{x}}-\hat{\boldsymbol{y}}-$polarized light is exclusively allowed at Y valley. As shown in Fig. 3(d), the overwhelming majority of optical transitions around the valleys preserve complete linear polarization selectivity.
@@
Just as the introduction of Berry curvature advanced our understanding of optical selection rules for circularly polarized light, our incorporation of the quantum metric also represents a fundamental advance for optical selection rules for linearly polarized light and enables new physics: 1. The direct experimental measurement of the quantum metric is a growing focus [43, 44], and quantum metric-based optical selection rules indicate that one can directly estimate the quantum metric by the absorption difference for orthogonal linear polarizations. 2. Due to the significant enhancement of the stiffness of exciton condensates by the quantum metric [45, 46], we anticipate the remarkable exciton condensates under quantum metric-based optical selection. 3. The quantum metric-based optical selection rules make widely investigated AMs a good playground for optical selection rules, since a significant number of reported abundant AMs exhibit mirror or mirror-like symmetry [47]. 4. The quantum metric-based optical selection rules unlock the potential of the $\Gamma$ valley without circular polarization selection for application in optical selection rules.